# What is glacier sliding?

### v1

@@ 3.3 Roughness, form drag, and temperate ice
In polythermal glacier and ice-sheet settings, a layer of lower-viscosity basal temperate ice (Fig. 4j) may furthermore modulate the complex motion patterns caused by topographic roughness (Krabbendam,, 2016; Law and others,, 2023), with greater thicknesses of temperate ice found in topographic troughs than peaks, and the formation of a shear band at the top of the temperate zone (Law and others,, 2023; Liu and others,, 2023). The rheology of temperate ice, and in particular temperate ice with an impurity content, is very poorly constrained and possibly even linear viscous (Lliboutry,, 1971; Adams and others,, 2021), making future studies into temperate ice rheology an urgent objective.
@@
The influence of roughness on sliding can be theorised through form drag. Kyrke-Smith and others, (2018) consider the influence of increasing the spatial resolution of a numerical model from 5 km to 0.5 km (and therefore also increasing the fidelity of the underlying topography) on the basal traction obtained from an inversion procedure. Intuitively, Kyrke-Smith and others, (2018) find that as spatial resolution is decreased, basal traction increases. Form drag is then described as the area-averaged basal traction with high topographic fidelity subtracted from the area-averaged basal traction with low topographic fidelity, i.e. the basal traction arising from topographic obstacles not explicitly represented by the basal boundary position. This clearly aligns with the idea that form drag can be theorised as part of the inner-flow which collapses to the modelled bed surface (Section 2). At a much smaller scale, Minchew and Joughin, (2020) comment on Zoet and Iverson, (2020) and invoke form drag as the viscous resistance that arises as ice deforms over entrained clast ‘micro-topography’ in laboratory soft-bed sliding experiments before the sediment reaches its yield strength described previously in Section 3.1.
@@
Form drag may therefore be a very useful concept in descriptions of glacier sliding, but its use in glaciology differs fundamentally from form drag as originally described for aerodynamic applications. The differences are sufficiently large that it may be more appropriate to use a different phrase altogether but we maintain the the phrase ‘form drag’ for continuity with previous studies. In aerodynamic applications, form drag refers to the force acting on a solid body moving through a viscous and locally-turbulent fluid that is opposite and parallel to its velocity and which is not accounted for by induced drag, where the induced drag is a force resulting from turbulent vorticies connected to the lift on the body (Batchelor,, 1967, sections 5.11, 7.8). Form drag has also been used in atmospheric and ocean dynamics, where it typically refers to the entire drag force resulting from turbulence around an obstacle or obstacles fixed to the surface, rather than a division of it (e.g. Arya,, 1973; Renfrew and others,, 2019; Jagannathan and others,, 2023). Difficulties are still encountered when applying the concept of form drag to settings where there is a fixed continuous surface however, even if turbulence is present (MacCready and others,, 2003).
Under the inner-outer flow framework described in Section 2 we fully define form drag as the viscous creep contribution to the inner flow not accounted for by slip processes occurring directly at the ice-bed interface or within subglacial sediment. In this manner the resistance from form drag is
^$\hat{\mathcal{F}}=\tau_{b}-\hat{\mathcal{P}_{s_{1}}}-\hat{\mathcal{P}_{s_{1}}}-...-\hat{\mathcal{P}_{s_{n}}}$ (7)
where $\hat{\mathcal{P}_{s}}$ terms are traction provided by sediment or slip processes and which can be expanded to an area $R$ in the same manner as Eq. 5. The contribution of form drag to sliding velocity then follows Eqs. 2 and 4. We consider temperate ice processes (this Section) and basal ice layer processes (Section 3.6) within the inner flow to be components of form drag such that
^$\hat{\mathcal{F}}=\hat{\mathcal{P}_{d_{1}}}+\hat{\mathcal{P}_{d_{2}}}+...+\hat{\mathcal{P}_{d_{n}}}$ (8)
where $\hat{\mathcal{P}_{d}}$ terms refer to sliding sub-processes occurring within the ice itself within the inner flow. Eq. 8 can also be expanded to an area $R$ or to describe the sliding velocity component in the established way.
3.4 Stick-slip and continuous sliding
Basal icequakes occur due to a rapid release of elastically stored energy, resulting in ice motion that is transiently far higher than the long-term average (Winberry and others,, 2009; Podolskiy and Walter,, 2016). Moment magnitudes have been recorded from negative in alpine settings (Helmstetter and others,, 2015) to magnitude seven at Whillans Ice Stream in West Antarctica (Wiens and others,, 2008). For a basal icequake to occur, the local frictional shear strength of the bed must be lower when the sliding interface is in motion than when it is static, i.e. $\mu_{t}$ must be lower than $\mu_{s}$ (Bahr and Rundle,, 1996; Rice and others,, 2001). Stick-slip is usually associated with diurnal or tidal variations in subglacial water pressure allowing a clear build-up period for elastic energy (Bahr and Rundle,, 1996; Bindschadler and others,, 2003; Walter and others,, 2008; Stevens and others,, 2024) but basal icequakes are also recorded outside of such cycles (Hubbard,, 2002). Stick-slip behaviour is usually described through a rate-and-state framework (Rice and others,, 2001) – as often applied in geological-fault settings (Gomberg and others,, 2000; Appendix D) – where a state variable, $\psi$, is introduced to account for the strength of the fault (Rice and others,, 2001; van den Ende and others,, 2018), usually based on the slip displacement (Ruina,, 1983) or its time-dependent evolution (Dieterich,, 1979) giving
$\tau_{b}=f(u_{b},N,\psi)\>.$ (9)
In contrast to Eq. 1, the expectation is that $u_{b}$ in Eq. 9 can vary rapidly in both time and space. A rate-and-state friction relationship can still account for the aseismic steady slip of a fault (van den Ende and others,, 2018) but is not well-suited for models entirely excluding rapid velocity changes and elastic behaviour by design.
Quantifying the contribution of stick-slip basal sliding to glacier motion is challenging due to large uncertainties in fault dimensions, slip displacement, and stress calculated from seismic data (Abercrombie,, 2015). Some estimates suggest stick-slip accounts for all basal sliding over an annual period (Helmstetter and others,, 2015), while in other cases, coseismic and aseismic regions coexist (Barcheck and others,, 2020; Kufner and others,, 2021) – analogous to coexisting seismogenic and creeping faults in actively deforming geological faults (e.g. Azzaro and others,, 2020).
Despite the growing interest in stick-slip motion in glaciers (e.g. Podolskiy and Walter,, 2016; Zoet and others,, 2020) there are limited connections between rate-and-state theory (of form Eq. 9) and continuous basal traction theory relationships (of form Eq. 1). We are only aware of the statistical mechanical approach of Bahr and Rundle, (1995), where the glacier sole is represented as numerous elastically interconnected blocks with stochastically varying roughness parameters and the similar but simplified 1D approach of (Köpfli and others,, 2022). Bahr and Rundle’s approach, intended for thin, steep, valley glaciers, finds that the average total force on an individual block (including elastic forces from adjacent blocks) is proportional to the area-average basal traction, $\tau_{b}$. If this holds for greater ice thicknesses and larger scales then stick-slip behaviour alone will not invalidate continuous basal traction relationships over a sufficiently large grid cell, but much further work is required to integrate this theory with recent developments in understanding the stick-slip behaviour of glaciers, and how this can relate to the spatial scales of numerical model grid cell resolution.
3.5 Overlap between soft- and hard-bed sliding
Most glacier and ice sheet modelling studies use basal traction relationships that implicitly assume the bed is either soft or hard across the entire domain with the ostensibly hard-bed Weertman-type sliding relationship (or equivalently, pseudo-plastic sliding with $2\geq m\geq 4$) being the modal choice in ISMIP6 experiments. In some situations a hard-soft division is appropriate, for example most studies are in agreement that extensive, relatively planar sediment lies beneath the ice streams draining into the Ronne and Ross ice shelves in Antarctica (Tulaczyk and others,, 2000; Vaughan and others,, 2003; Peters and others,, 2007), while geographically limited areas of Canadian continental shield may qualify as ‘true’, relatively planar, sediment-free hard beds (Slaymaker and Kovanen,, 2017). Many individual valley glaciers can also be reasonably categorised as predominantly soft- (e.g. Murray and Porter,, 2001) or hard-bedded (e.g. Hubbard,, 2002) (Section 6.1). However, taking Isunnguata Sermia in west Greenland as one well-studied example, borehole data mostly from topographic highs are used to suggest hard-bed conditions are dominant (Harper and others,, 2017; Maier and others,, 2019) while nearby seismic surveys are used to suggest the opposite (Booth and others,, 2012; Dow and others,, 2013; Kulessa and others,, 2017). Meanwhile, ice-marginal studies in west Greenland (Klint and others,, 2013) find a complex mix of bedrock, sediment-filled depressions, and extensive sandurs filling valley bottoms in front of the major land-terminating outlets (Grocott and McCaffrey,, 2017). In the case of the seismic survey of Kulessa and others, (2017) the recorded seismograms also allow for hard-bed conditions at topographic highs which would be in agreement with Harper and others, (2017). Similarly, recent work at Thwaites Glacier in Antarctica further suggests the coexistence of soft- and hard-bed regions over distances less than 10 km (Jordan and others,, 2023), supported by distinct along-flow variations in roughness (Holschuh and others,, 2020). While soft- or hard-bed categories can be a reasonable distinction for individual glaciers, all available evidence indicates that a more appropriate default for present and palaeo ice-sheets is a mixed continuum between the two end-members.
Some recent attention has been focused on the interaction between soft- and hard-beds, with Koellner and others, (2019) considering 2D flow over sinusoidal beds with wavelengths from 8-60 km and setting the value of $m$ in the Weertman-type sliding relationship (Table 1) to 8 to approximate regularised-Coulomb sliding in sinusoid troughs, while keeping it at 3 over the sinusoid peaks. This demonstrates that a hybrid $m=3\leftrightarrow 8$ Weertman rheology results in a modelled glacier response between $m=3$ and $m=8$ end-members. Given sliding rates at the modelled ice-bed interface can vary significantly over much shorter distances (Law and others,, 2023) than considered in Koellner and others, (2019), further work is needed to effectively isolate the influence of realistic variations in soft- and hard-bed sliding, and to assess whether hard-bed highs and soft-bed depressions are a reasonable model for the geologic composition of the Greenland and Antarctic ice sheets. Joughin and others, (2010) also apply a plastic sliding relationship to inferred soft-bedded regions and Weertman-type sliding with $m=3$ to inferred hard-bedded regions Joughin and others, (2009) at Pine Island Glacier, Antarctica, finding good agreement in this mixed model over an 8-year period (Section 5).
The recent suggestion that similar forms of the empirically derived sliding relationships for hard-bed cavitation (Schoof,, 2005; Gagliardini and others,, 2007) and soft-bed sediment deformation present a ‘universal glacier slip law’ (Minchew and Joughin,, 2020), excluding the other sliding sub-processes discussed in this paper merits reconsideration. It is challenging to reconcile the relationships of Zoet and Iverson, (2020) and Helanow and others, (2021) without requiring an unrealistically low friction angle in Zoet and Iverson, (2020) (Appendix A), indicating that the set-up of Zoet and Iverson, (2020) produces a stronger bed than Helanow and others, (2021) under equivalent effective pressure and velocities. Despite the general conception that soft beds are weaker than their hard-bed equivalents (e.g. Koellner and others,, 2019; Joughin and others,, 2009), the strength difference may be explained by (i) a natural requirement for lower subglacial water pressure in hard beds (by around 4%), which is supported by available evidence (e.g. Engelhardt and others,, 1990; Doyle and others,, 2018), (ii) non-negligible ice-rock interface friction, which is not included in Helanow and others, (2021), or (iii) unaccounted for form-drag occurring in most hard-bed settings that leads to a view of a stronger bed. There is some ground to cover yet to reconcile these differences, which may still not necessarily constitute a truly unified sliding relationship incorporating all sliding sub-processes (Section 6.4).
Last, Tsai and others, (2021) take an agnostic approach to the soft- and hard- bed division, focusing instead on the region of the bed where the hydrological system is expected to be active (where $\tau_{b}$ is locally set to 0) and inactive (where $\tau_{b}$ is locally determined using Weertman-type sliding). Under this formulation Tsai and others, (2021) find a relationship similar to Budd and others, (1984) sliding (their Eq. 6) which matches well with annual variations in velocity and moulin discharge at a land-terminating sector of the Greenland Ice Sheet. A simplification of this relationship may be well-applicable in large-scale models (Section 6.4).
@@ 3.6 The basal ice layer and ice rheology
Shear-ring laboratory experiments for the deformation of a 1-2 cm thick ice-sediment melange between thawed sediments and sediment (Hansen and others,, 2024) suggest similar regularised-Coulomb behaviour to shear-ring experiments of clean ice over sediment (Zoet and Iverson,, 2020). Meyer and others, (2018) also construct an analytical model arguing that ice infiltration into previously thawed sediments limits bed strength (in effect, plastic sliding) in pervasively soft-bedded regions. Beyond these studies however, we are not aware of experiments considering basal ice sequences exceeding a few cm in thickness or the large diversity of facies found in the basal ice layer and their interaction (Hubbard and others,, 2009).
@@
Four further rheological factors present significant complications for ice deformation in the vicinity of the bed and are hence captured at least partially within the inner flow. First, pre-Holocene ice with a higher dust concentration and typically smaller grain size deposited during the last glacial period deforms at a rate around 2.5 times that of Holocene ice under the same stress and temperature conditions (Paterson,, 1991). Such ice is widely present in the strata of the Greenland and Antarctic Ice Sheets (Macgregor and others,, 2015; Winter and others,, 2019; Ashmore and others,, 2020) but uncommon in glaciers and ice caps where it is limited to very slow moving areas (Thompson and others,, 1997), though valley glaciers generally feature much higher rates of englacial debris impurities (e.g. Goodsell and others,, 2005). Second, ice at depth has a highly anisotropic rheology due to the development of a strong crystallographic preferred orientation under consistent uni-directional deformation (Lile,, 1978; Baker,, 1981; Duval,, 1981; Wilson and Sim,, 2002). Such mechanical anisotropy presents major challenges in implementation, bench-marking, and interpretation within numerical modelling (Martín and others,, 2004; Gillet-Chaulet and others,, 2005) but is nonetheless likely of significant importance within viscous deformation of ice in the inner flow where strain rates are both high and spatially complex (Law and others,, 2023). To our knowledge, this aspect of glacier sliding is not covered by any existing glaciological literature. Third, folds close to the bed may arise from mechanical differences in ice layers (Whillans and Jezek,, 1987; Bons and others,, 2016; Zhang and others,, 2023) with the possibility to significantly influence the bulk rheological behaviour of the inner flow. Fourth, while $n=3$ is the default in glacier and ice-sheet models (in fact already rounded down from the 3.2 reported in Glen,, 1955), there are numerous field-observation based studies suggesting that $n=4$, representing a dislocation creep regime, is more appropriate in higher stress settings (away from ice-sheet ice divides, for example) (Bons and others,, 2018; Millstein and others,, 2022; Ranganathan and Minchew,, 2024) and perhaps also at areas dominated by pure shear (Gillet-Chaulet and others,, 2011). Considering a smooth BedMachine lower boundary in northern Greenland, $n=4$ dramatically decreases the area over which basal sliding is expected to contribute significantly to total surface displacement (Bons and others,, 2018). The influence of $n=4$ on form drag and sliding parameterisations, and its interaction with rough topography is to our knowledge entirely unexplored.

### v3

@@ 3.3 Roughness, form drag, and temperate ice
In polythermal glacier and ice-sheet settings, a layer of lower-viscosity basal temperate ice (Fig. 4j) may furthermore modulate the complex motion patterns caused by topographic roughness (Krabbendam,, 2016; Law and others,, 2023), with greater thicknesses of temperate ice found in topographic troughs than peaks, and the formation of a shear band at the top of the temperate zone (Law and others,, 2023; Liu and others,, 2023). The rheology of pure temperate ice has recently been constrained as likely linear viscous (Lliboutry,, 1971; Adams and others,, 2021; Schohn and others,, 2025), though the rheology of temperate ice with a high impurity content remains unconstrained through laboratory experiments to our knowledge.
@@
The influence of roughness on sliding can be theorised through form drag, even if it is at fist glance complicated to delineate the transition between ice deformation contributing to presently-derived hard bed sliding relationships (Section 3.2), and ice deformation contributing to form drag. Kyrke-Smith and others, (2018) consider the influence of increasing the spatial resolution of a numerical model from 5 km to 0.5 km (and therefore also increasing the fidelity of the underlying topography) on the basal traction obtained from an inversion procedure. Intuitively, Kyrke-Smith and others, (2018) find that as spatial resolution is decreased, basal traction increases. Form drag is then described as the area-averaged basal traction with high topographic fidelity subtracted from the area-averaged basal traction with low topographic fidelity, i.e. the basal traction arising from topographic obstacles not explicitly represented by the basal boundary position. This clearly aligns with the idea that form drag can be theorised as part of the inner-flow which collapses to the modelled bed surface (Section 2). At a much smaller scale, Minchew and Joughin, (2020b) comment on Zoet and Iverson, (2020) and invoke form drag as the viscous resistance that arises as ice deforms over entrained clast ‘micro-topography’ in laboratory soft-bed sliding experiments before the sediment reaches its yield strength described previously in Section 3.1.
@@
Form drag may therefore be a very useful concept in descriptions of glacier sliding, but its use in glaciology differs fundamentally from form drag as originally described for many aerodynamic applications. The differences are sufficiently large that it may be more appropriate to use a different phrase altogether but we maintain the the phrase ‘form drag’ for continuity with previous studies. In aerodynamic applications, form drag refers to the force acting on a solid body moving through a viscous and locally-turbulent fluid that is opposite and parallel to its velocity and which is not accounted for by induced drag, where the induced drag is a force resulting from turbulent vorticies connected to the lift on the body (Batchelor,, 1967, sections 5.11, 7.8). Form drag has also been used in atmospheric and ocean dynamics, where it typically refers to the entire drag force resulting from turbulence around an obstacle or obstacles fixed to the surface, rather than a division of it (e.g. Arya,, 1973; Renfrew and others,, 2019; Jagannathan and others,, 2023). Difficulties are still encountered when applying the concept of form drag to settings where there is a fixed continuous surface however, even if turbulence is present (MacCready and others,, 2003).
To circumvent the problem of defining form drag only at a specific scale – i.e. for ‘topographic’ roughness (Kyrke-Smith and others,, 2018) or for ‘sedimentary-clast’ roughness (Minchew and Joughin,, 2020a) – we define form drag here as $\boldsymbol{\mathcal{D}}_{T}+\boldsymbol{\mathcal{G}}$: the sum of all distributed viscous deformation and geometric contributions to the resistive stress operating over a given region within the inner layer. This means deformation processes across a very wide spatial scale (Fig. 2) including temperate ice (this section), basal ice layer processes (Section 3.4), and ’enhanced creep’ described in the original Weertman, (1957) paper are all taken as form drag components under our definition. We make this grouping lacking a consistent rule to meaningfully make distinctions between these processes.
@@ 3.4 The basal ice layer and ice rheology
Shear-ring laboratory experiments for the deformation of a 1-2 cm thick ice-sediment melange between thawed sediments and sediment (Hansen and others,, 2024) suggest similar regularised-Coulomb behaviour to shear-ring experiments of clean ice over sediment (Zoet and Iverson,, 2020). Meyer and others, (2018) also construct an analytical model arguing that ice infiltration into previously thawed sediments limits bed strength (in effect, plastic sliding) in pervasively soft-bedded regions. Beyond these studies however, we are not aware of experiments considering basal ice sequences exceeding a few cm in thickness or the large diversity of facies found in the basal ice layer (Hubbard and others,, 2009).
@@
Four further rheological factors present significant complications for ice deformation in the vicinity of the bed and are hence captured at least partially within the inner flow. First, pre-Holocene ice with a higher dust concentration and typically smaller grain size deposited during the last glacial period deforms at a rate around 2.5 times that of Holocene ice under the same stress and temperature conditions (Paterson,, 1991). Such ice is widely present in the strata of the Greenland and Antarctic Ice Sheets (Macgregor and others,, 2015; Winter and others,, 2019; Ashmore and others,, 2020) but uncommon in glaciers and ice caps where it is limited to very slow moving areas (Thompson and others,, 1997), though valley glaciers generally feature much higher rates of englacial debris impurities in any case (e.g. Goodsell and others,, 2005). Second, ice at depth has a highly anisotropic rheology due to the development of a strong crystallographic preferred orientation under consistent uni-directional deformation (Lile,, 1978; Baker,, 1981; Duval,, 1981; Wilson and Sim,, 2002). Such mechanical anisotropy presents major challenges in implementation, bench-marking, and interpretation within numerical modelling (Martín and others,, 2004; Gillet-Chaulet and others,, 2005) but is nonetheless likely of significant importance within viscous deformation of ice in the inner flow where strain rates are both high and spatially complex (Law and others,, 2023). To our knowledge, this aspect of glacier sliding is not covered by any existing glaciological literature. Third, folds close to the bed may arise from mechanical differences in ice layers (Whillans and Jezek,, 1987; Bons and others,, 2016; Zhang and others,, 2023) with the possibility to significantly influence the bulk rheological behaviour of the inner flow. Fourth, while $n=3$ is the default in glacier and ice-sheet models (in fact already rounded down from the 3.2 reported in Glen,, 1955), there are numerous field-observation based studies suggesting that for cold (i.e. not temperate) ice $n=4$, representing a dislocation creep regime, is more appropriate in higher stress settings (away from ice-sheet ice divides, for example) (Bons and others,, 2018; Millstein and others,, 2022; Ranganathan and Minchew,, 2024) and perhaps also at areas dominated by pure shear (Gillet-Chaulet and others,, 2011). Considering a smooth BedMachine lower boundary in northern Greenland, $n=4$ dramatically decreases the area over which basal sliding is expected to contribute significantly to total surface displacement (Bons and others,, 2018). The influence of $n=4$ on form drag and sliding parameterisations, and its interaction with rough topography is to our knowledge entirely unexplored.
@@
3.5 Stick-slip and continuous sliding
Basal icequakes occur due to a rapid release of elastically stored energy, resulting in ice motion that is transiently far higher than the long-term average (Winberry and others,, 2009; Podolskiy and Walter,, 2016). Moment magnitudes have been recorded from negative in alpine settings (Helmstetter and others,, 2015) to magnitude seven at Whillans Ice Stream in West Antarctica (Wiens and others,, 2008). For a basal icequake to occur, the local frictional shear strength of the bed must be lower when the sliding interface is in motion than when it is static, i.e. $\mu_{t}$ must be lower than $\mu_{s}$ (Bahr and Rundle,, 1996; Rice and others,, 2001) describing slip-weakening behaviour. Stick-slip is usually associated with diurnal or tidal variations in subglacial water pressure allowing a clear build-up period for elastic energy (Bahr and Rundle,, 1996; Bindschadler and others,, 2003; Walter and others,, 2008; Stevens and others,, 2024) but basal icequakes are also recorded outside of such cycles (Hubbard,, 2002). Stick-slip behaviour is usually described through a rate-and-state framework (Rice and others,, 2001) – as often applied in geological-fault settings (Gomberg and others,, 2000; Appendix A4) – where a state variable, $\psi$, is introduced to account for the strength of the fault (Rice and others,, 2001; van den Ende and others,, 2018), usually based on the slip displacement (Ruina,, 1983) or its time-dependent evolution (Dieterich,, 1979) giving
$\tau_{b}=f(u_{b},N,\psi)\>.$ (15)
In contrast to Eq. 1, the expectation is that $u_{b}$ in Eq. 15 can vary rapidly in both time and space. A rate-and-state friction relationship can still account for the aseismic rate-strengthening steady slip of a fault (van den Ende and others,, 2018) but is not well-suited for models entirely excluding rapid velocity changes and elastic behaviour by design.
Quantifying the contribution of stick-slip basal sliding to glacier motion is challenging due to large uncertainties in fault dimensions, slip displacement, and stress calculated from seismic data (Abercrombie,, 2015). Some estimates suggest stick-slip displacement accounts for all basal sliding over an annual period (Helmstetter and others,, 2015), while in other cases, coseismic and aseismic regions coexist (Barcheck and others,, 2020; Kufner and others,, 2021) – analogous to coexisting seismogenic and creeping faults in actively deforming geological faults (e.g. Azzaro and others,, 2020).
Despite the growing interest in stick-slip motion in glaciers (e.g. Podolskiy and Walter,, 2016; Zoet and others,, 2020) there are limited connections between rate-and-state theory (of form Eq. 15) and continuous basal traction theory relationships (of form Eq. 1). We are only aware of the statistical mechanical approach of Bahr and Rundle, (1995), where the glacier sole is represented as numerous elastically interconnected blocks with stochastically varying roughness parameters and the similar but simplified 1D approach of (Köpfli and others,, 2022). Bahr and Rundle’s approach, intended for thin, steep, valley glaciers, finds that the average total force on an individual block (including elastic forces from adjacent blocks) is proportional to the area-average basal traction, $\tau_{b}$. If this holds for greater ice thicknesses and larger scales then stick-slip behaviour alone will not invalidate continuous basal traction relationships over a sufficiently large grid cell (or representative surface area), but much further work is required to integrate this theory with recent developments in understanding the stick-slip behaviour of glaciers, and how this can relate to the spatial scales of numerical model grid cell resolution.
3.6 Overlap between soft- and hard-bed sliding
Most glacier and ice sheet modelling studies use basal traction relationships that implicitly assume the bed is either soft or hard across the entire domain with the ostensibly hard-bed Weertman-type sliding relationship (or equivalently, power-law/pseudo-plastic sliding with $2\geq m\geq 4$) being the modal choice in ISMIP6 experiments. In some situations a hard-soft division is appropriate, for example most studies are in agreement that extensive, relatively planar sediment lies beneath the ice streams draining into the Ronne and Ross ice shelves in Antarctica (Tulaczyk and others,, 2000; Vaughan and others,, 2003; Peters and others,, 2007), while geographically limited areas of Canadian continental shield may qualify as ‘true’, relatively planar, sediment-free hard beds (Slaymaker and Kovanen,, 2017). Many individual valley glaciers can also be reasonably categorised as predominantly soft- (e.g. Murray and Porter,, 2001) or hard-bedded (e.g. Hubbard,, 2002) (Section LABEL:s:glaciericesheet). However, taking Isunnguata Sermia in west Greenland as one well-studied example, borehole data mostly from topographic highs are used to suggest hard-bed conditions are dominant (Harper and others,, 2017; Maier and others,, 2019) while nearby seismic surveys are used to suggest the opposite (Booth and others,, 2012; Dow and others,, 2013; Kulessa and others,, 2017). Meanwhile, ice-marginal studies in west Greenland (Klint and others,, 2013) find a complex mix of bedrock, sediment-filled depressions, and extensive sandurs filling valley bottoms in front of the major land-terminating outlets (Grocott and McCaffrey,, 2017). In the case of the seismic survey of Kulessa and others, (2017) the recorded seismograms also allow for hard-bed conditions at topographic highs which would be in agreement with Harper and others, (2017). Similarly, recent work at Thwaites Glacier in Antarctica further suggests the coexistence of soft- and hard-bed regions over distances less than 10 km (Jordan and others,, 2023), supported by distinct along-flow variations in roughness (Holschuh and others,, 2020). While soft- or hard-bed categories can be a reasonable distinction for individual glaciers, all available evidence indicates that a more appropriate default for present and palaeo ice-sheets at large is a mixed bed condition vacillating between the two end-members.
Some recent attention has been focused on the interaction between soft- and hard-beds, with Koellner and others, (2019) considering 2D flow over sinusoidal beds with wavelengths from 8-60 km and setting the value of $m$ in the power-law sliding relationship (Table 1) to 8 to approximate regularised-Coulomb sliding in sinusoid troughs, while keeping it at 3 over the sinusoid peaks. This demonstrates that a hybrid $m=3\leftrightarrow 8$ Weertman rheology results in a modelled glacier response between $m=3$ and $m=8$ end-members. Given sliding rates at the modelled ice-bed interface can vary significantly over much shorter distances (Law and others,, 2023) than considered in Koellner and others, (2019), further work is needed to effectively isolate the influence of realistic variations in soft- and hard-bed sliding, and to assess whether hard-bed highs and soft-bed depressions are a reasonable model for the geologic composition of the Greenland and Antarctic ice sheets. Joughin and others, (2010) also apply a plastic sliding relationship to inferred soft-bedded regions and power-law sliding with $m=3$ to inferred hard-bedded regions Joughin and others, (2009) at Pine Island Glacier, Antarctica, finding good agreement in this mixed model over an 8-year period (Section 5).
From empirical studies at small (<25 m) scales, both the derived sliding relationship for hard-bed cavitation (Schoof,, 2005; Gagliardini and others,, 2007; Helanow and others,, 2021) and slip over sediments (Zoet and Iverson,, 2020) can yield a bounded traction relationship, leading to suggestions that regularised-Coulomb then represents a ‘universal glacier slip law’ for both hard and soft beds (Minchew and Joughin,, 2020b). We cover large-scale remote-sensing/modelling studies which indicate this may not be the case in Section 5, and consider the applicability of a ‘unified’ sliding relationship in Section 6.3, but focus briefly here on the differences that still remain between the regularised-Coulomb sliding relationships of Helanow and others, (2021) and Zoet and Iverson, (2020). In order to match the relationships of Zoet and Iverson, (2020) and Helanow and others, (2021) an unrealistically low friction angle is required in the equation of Zoet and Iverson, (2020) (Fig. A1), indicating that the set-up of Zoet and Iverson, (2020) produces a stronger bed than Helanow and others, (2021) under equivalent effective pressure and velocities, despite the general conception that soft beds are weaker than their hard-bed equivalents (e.g. Koellner and others,, 2019; Joughin and others,, 2009). It is possible that the strength difference is explained by (i) a natural requirement for lower subglacial water pressure in hard beds (by around 4%), which is supported by available evidence (e.g. Engelhardt and others,, 1990; Doyle and others,, 2018), (ii) non-negligible ice-rock interface friction, which is not included in Helanow and others, (2021), or (iii) unaccounted for form drag (i.e. an increase in $\boldsymbol{\mathcal{D}}_{T}$ within Eq. 8 if sliding with cavitation described in Helanow and others,, 2021, is taken as a process within $\boldsymbol{\mathcal{S}}_{T}$) occurring in many hard-bed settings which results in a view of a stronger bed.
Last, Tsai and others, (2021) take an agnostic approach to the soft- and hard- bed division, focusing instead on the region of the bed where the hydrological system is expected to be active (where $\tau_{b}$ is locally set to 0) and inactive (where $\tau_{b}$ is locally determined using power-law sliding). Under this formulation Tsai and others, (2021) find a relationship similar to Budd and others, (1984) sliding (their Eq. 6) which matches well with annual variations in velocity and moulin discharge at a land-terminating sector of the Greenland Ice Sheet. This relationship may be well-applicable in large-scale models (Section 6.3).